# Joint Resource-Power Allocation and UE Rank Selection in Multi-User MIMO Systems with Linear Transceivers

### v1

@@ Abstract
Next-generation wireless networks aim to deliver data speeds much faster than 5G. This requires base stations with lots of antennas and a large operating bandwidth. These advanced base stations are expected to serve several multi-antenna user-equipment (UEs) simultaneously on the same time-frequency resources on both the uplink and the downlink. The UE data rates are affected by the following three main factors: UE rank, which refers to the number of data layers used by each UE, UE frequency allocation, which refers to the assignment of slices of the overall frequency band to use for each UE in an orthogonal frequency-division multiplexing (OFDM) system, and UE power allocation/control, which refers to the power allocated by the base station for data transmission to each UE on the downlink or the power used by the UE to send data on the uplink. Since multiple UEs are to be simultaneously served, the type of precoder used for downlink transmission and the type of receiver used for uplink reception predominantly influence these three aforementioned factors and the resulting overall UE throughput. This paper addresses the problem of jointly selecting these three parameters specifically when zero-forcing (ZF) precoders are used for downlink transmission and linear minimum mean square error (LMMSE) receivers are employed for uplink reception.
@@ I Introduction
In single-user MIMO (SU-MIMO) transmissions, the orthogonal frequency-division multiple access (OFDMA) scheme avoids inter-user interference within the same cell. However, when multiple UEs are co-scheduled for transmission in the same time-frequency resources in a time slot, their respective transmitted data signals from the BS (downlink) or their respective received data signals at the BS (uplink) could potentially interfere with one another. The exact degree of this interference depends on the spatial correlation between channels of the paired UEs, the type of MU-MIMO precoder used at the BS (downlink), and the type of MU-MIMO detector used at the BS (uplink). Therefore, the resulting UE data rates depend on how the frequency resources are shared between the UEs and this in turn influences the respective UE ranks and power allocation. Much of the literature has focused on optimizing these three procedures separately, but this leads to numerous balloon-effects, where optimizing one resource reduces the gains achievable by optimizing another. Furthermore, up until the recent advent of 5G NR, MU-MIMO was not commonly used, which has also delayed the development of joint solutions for this problem. Practical solutions for the joint allocation of time-frequency resources, power control and rank selection remain elusive and are much needed to avoid inter-user interference and for 6G MU-MIMO to succeed.
This paper addresses the problem of joint resource-power allocation and rank selection for a class of linear transceivers. In particular, we consider zero-forcing (ZF) precoders [1] for downlink transmission and linear minimum mean square error (LMMSE) receivers for uplink reception. We show that both these problems can be formulated under a common framework. Since the problems involve mixed integer programming and finding the optimal solution could be computationally intensive, we provide a practically-feasible approach to obtain a reasonably good (even if possibly sub-optimal) solution. To the best of our knowledge, this is the first such work on jointly selecting the three aforementioned features based on the type of precoder or receiver used.
@@
Notation: Boldface upper-case (lower-case) letters denote matrices (vectors). The field of complex numbers and the field of real numbers are respectively denoted by $\mathbb{C}$ and $\mathbb{R}$. For any set ${\cal S}$, $|{\cal S}|$ denotes its cardinality if it is finite and the notation $\mathbf{X}\in{\cal S}^{m\times n}$ denotes that $\mathbf{X}$ is a matrix of size $m\times n$ with each entry taking values from a set ${\cal S}$. For a matrix $\mathbf{X}$, its transpose and Hermitian transpose are respectively denoted by $\mathbf{X}^{T}$ and $\mathbf{X}^{H}$, Frobenius norm by $\|\mathbf{X}\|$, and its $(i,j)^{th}$ entry by $[\mathbf{X}]_{i,j}$. The block-diagonal matrix with diagonal blocks $\mathbf{D}_{1},\mathbf{D}_{2},\cdots,\mathbf{D}_{n}$ is denoted by $\mathrm{diag}\left[\mathbf{D}_{1},\mathbf{D}_{2},\cdots,\mathbf{D}_{n}\right]$, and the same notation is used for denoting a diagonal matrix with diagonal elements $d_{1},\cdots,d_{n}$. The identity matrix and the null matrix are respectively denoted by $\mathbf{I}$ and $\mathbf{O}$ with their sizes understood from context. Finally, $\mathbb{I}_{\{A\}}$ denotes the indicator function for event $A$; $\mathbb{I}_{\{A\}}=1$ if $A$ occurs, and is 00 otherwise.
@@ II System Model
We consider an OFDM-based multiple-input multiple-output (MIMO) communication system with $n_{F}$ subcarriers. A resource element (RE) corresponds to a subcarrier-OFDM symbol pair in the OFDM grid. A slot consists of $T$ OFDM symbols (usually $T=12$ or $14$) and a PRB consists of 12 subcarriers. Suppose that the BS is equipped with $n_{B}$ antennas and serves a set of $N_{UE}\geq 1$ co-scheduled UE, with UE $i$ equipped with $n_{U}^{(i)}$ antennas, $i=1,\cdots,N_{UE}$. Each UE $i$ is assigned a rank (also called the number of data layers or spatial streams assigned for that UE) $n_{l}^{(i)}\leq n_{U}^{(i)}$ with $n_{L}\triangleq\sum_{i=1}^{N_{UE}}n_{l}^{(i)}\leq n_{B}$. Let $\mathbf{H}_{i,f,t}\in\mathbb{C}^{n_{B}\times n_{U}^{(i)}}$ denote the channel matrix between UE $i$ and the BS for RE $(f,t)$, where $f$ corresponds to the subcarrier index and $t=1,\cdots,T$, to the OFDM symbol index. While it is assumed that the BS has perfect knowledge of this channel-state information (CSI), in practice, only imperfect estimates are available, either through sounding reference signal (SRS) [8, Sec. 6.4.1.4] on the uplink in a time division duplex (TDD) system or through UE CSI feedback in a frequency division duplex (FDD) system. Moreover, these estimates are only available at a certain granularity (e.g., one channel estimate for every two consecutive PRBs). In this paper, we call this set of contiguous PRBs for which a single channel estimate is available a physical resource block group (PRBG). This can also be a group of PRBs over which the channel is approximately constant. Therefore, resource allocation refers to the assignment of PRBGs to the $N_{UE}$ UEs that are to be served in the slot in context in order to optimize a certain performance metric, with possibly multiple UEs sharing the same PRBG. It is assumed that once a PRBG is assigned to a UE, all the $T$ OFDM symbols are used for communication by the UE in that PRBG. The constellation used for communication between the BS and UE $i$ is a unit energy $2^{m_{i}}$-QAM constellation of size $2^{m_{i}}$ for some $m_{i}=2,4,6,8,\cdots$, and is denoted by ${\cal Q}_{i}$.
@@ II-B Downlink transmission
Let ${\cal I}_{f,t}\subseteq\{1,\cdots,N_{UE}\}$ denote the set of UEs that have been allocated the RE $(f,t)$, and let $n_{L,f,t}\triangleq\sum_{i\in{\cal I}_{f,t}}n_{l}^{(i)}$ be the total number of active layers in the RE $(f,t)$. We denote by $\mathbf{W}_{i,f,t}\in\mathbb{C}^{n_{B}\times n_{l}^{(i)}}$ the part of the precoding matrix associated with UE $i$ if $i\in{\cal I}_{f,t}$, and the overall precoding matrix $\mathbf{W}_{f,t}\in\mathbb{C}^{n_{B}\times n_{L,f,t}}$ used by the BS is obtained by stacking the matrices $\{\mathbf{W}_{i,f,t},i\in{\cal I}_{f,t}\}$ column-wise, with each column of $\mathbf{W}_{f,t}$ having unit norm. Further, let $a_{k,i,j,f,t}\triangleq|\left[\mathbf{W}_{i,f,t}\right]_{k,j}|^{2}$, $k=1\cdots,n_{B},j=1,\cdots,n_{l}^{(i)}$. The BS transmits $\mathbf{x}_{i,f,t}\triangleq\left[x_{1,i,f,t},\cdots,x_{n_{l}^{(i)},i,f,t}\right]^{T}\in{\cal Q}_{i}^{n_{l}^{(i)}\times 1}$ on RE $(f,t)$ to UE $i$ if $i\in{\cal I}_{f,t}$, and uses a transmit power $p_{i,j,f,t}>0$ for the $j^{th}$ layer of UE $i$ ($p_{i,j,f,t}=0$ if $i\notin{\cal I}_{f,t}$), and let $\mathbf{D}_{i,f,t}\triangleq\mathrm{diag}\left[\sqrt{p_{i,1,f,t}},\cdots,\sqrt{p_{i,n_{l}^{(i)},f,t}}\right]$. With a maximum BS transmit power of $P_{B,max}$ and a per-antenna power constraint (PAPC) of $P_{ant}=P_{B,max}/n_{B}$, it follows that
$\sum_{k=1}^{n_{B}}\sum_{i\in{\cal I}_{f,t}}\sum_{j=1}^{n_{l}^{(i)}}\sum_{f=1}^{n_{F}}a_{k,i,j,f,t}p_{i,j,f,t}\leq P_{B,max},\forall t=1,\cdots,T,$ (6)
$\sum_{i\in{\cal I}_{f,t}}\sum_{j=1}^{n_{l}^{(i)}}\sum_{f=1}^{n_{F}}a_{k,i,j,f,t}p_{i,j,f,t}\leq P_{ant},\forall k=1,\cdots,n_{B},\forall t=1,\cdots,T.$ (7)
In this paper, we consider the class of precoders called ZF precoders [1] that satisfy, for any $j,i\in{\cal I}_{f,t}$, $\mathbf{H}_{j,f,t}^{T}\mathbf{W}_{i,f,t}=\mathbf{O}$ when $j\neq i$, and $\mathbf{H}_{i,f,t}^{T}\mathbf{W}_{i,f,t}=\mathbf{U}_{i,f,t}\mathbf{\Lambda}_{i,f,t}$ for some matrix $\mathbf{U}_{i,f,t}\in\mathbb{C}^{n_{U}^{(i)}\times n_{l}^{(i)}}$ satisfying $\mathbf{U}_{i,f,t}^{H}\mathbf{U}_{i,f,t}^{H}=\mathbf{I}$ and $\mathbf{\Lambda}_{i,f,t}=\mathrm{diag}\left[\sqrt{\lambda_{i,1,f,t}},\cdots,\sqrt{\lambda_{i,n_{l}^{(i)},f,t}}\right]$ with positive real-valued diagonal elements. With this, the signal model for the received signal at UE $i$ on RE $(f,t)$ is
$\mathbf{y}_{i,f,t}=\mathbf{H}_{i,f,t}\mathbf{W}_{i,f,t}\mathbf{D}_{i,f,t}\mathbf{x}_{i,f,t}+\mathbf{n}_{i,f,t}$ (8)
@@ II-C Expected UE Rates
Modern communication systems use bit-interleaved coded modulation (BICM) [10] where, due to the usage of scrambling and interleaving, the transmitted bits in an RE are approximately independent and identically distributed. Let ${\cal G}_{i}\triangleq\{(f,t)|i\in{\cal I}_{f,t}\}$ denote the indices of the RE allocated for communication between the BS and UE $i$. In practical communication systems, there is finite number of modulation and coding schemes (MCSs). For example, 5G NR Table 2 supports up to $27$ MCS levels [8, Table 5.1.3.1-2], with the lowest level using $4$-QAM and a rate-$\frac{120}{1024}$ low-density parity-check (LDPC) code (spectral efficiency (SE) given by $\mathop{\mathrm{log}_{2}}\left(4\right).\left(120/1024\right)=0.234$) and the highest level using $256$-QAM and a rate-$\frac{948}{1024}$ LDPC code (SE equals $\mathop{\mathrm{log}_{2}}\left(256\right).\left(948/1024\right)=7.40$). We denote the minimum MCS SE by $r_{min}$ and the maximum rate possible by $r_{max}=\mathop{\mathrm{log}_{2}}\left(|Q_{max}|\right)$, respectively, where $Q_{max}$ is the highest order constellation used in the system. The achievable rate $R_{UE,i}$ (in bits per slot) for UE $i$ is upper-bounded as
@@
where $\rho_{i,j,f,t,i}$ is the post-equalization SINR for the $j^{th}$ received symbol of UE $i$ in the RE indexed by $(f,t)$, and $I_{Q}(x)$ refers to the mutual-information (MI) between the input and the output in an AWGN channel for constellation $Q$ and an SNR of $x$. Since there is a minimum MCS SE $r_{min}$, reliable communication is not guaranteed if the expression in (11) is lower than $\left|{\cal G}_{i}\right|n_{l}^{(i)}r_{min}$. Therefore, the upper bound on the rate for UE $i$ is
$R_{UE,i,max}=\left|{\cal G}_{i}\right|n_{l}^{(i)}x\mathbb{I}_{\{x>r_{min}\}},~{}~{}~{}x=\frac{1}{n_{l}^{(i)}\left|{\cal G}_{i}\right|}\sum_{(f,t)\in{\cal G}_{i}}\sum_{j=1}^{n_{l}^{(i)}}I_{Q_{i}}\left(\rho_{i,j,f,t}\right).$ (12)

### v2

@@ Abstract
Next-generation wireless networks aim to deliver data speeds much faster than 5G. This requires base stations with lots of antennas and a large operating bandwidth. These advanced base stations are expected to serve several multi-antenna user-equipment (UEs) simultaneously on the same time-frequency resources on both the uplink and the downlink. The UE data rates are affected by the following three main factors: UE rank, which refers to the number of data layers used by each UE, UE frequency allocation, which refers to the assignment of slices of the overall frequency band to use for each UE in an orthogonal frequency-division multiplexing (OFDM) system, and UE power allocation/control, which refers to the allocation of power by the base station for data transmission to each UE on the downlink or the power used by each UE to send data on the uplink. Since multiple UEs are to be simultaneously served, the type of precoder used for downlink transmission and the type of receiver used for uplink reception predominantly influence these three aforementioned factors and the resulting overall UE throughput. This paper addresses the problem of jointly selecting these three parameters specifically when zero-forcing (ZF) precoders are used for downlink transmission and linear minimum mean square error (LMMSE) receivers are employed for uplink reception.
@@ I Introduction
In single-user MIMO (SU-MIMO) transmissions, the orthogonal frequency-division multiple access (OFDMA) scheme avoids inter-user interference within the same cell. However, when multiple UEs are co-scheduled for transmission in the same time-frequency resources in a time slot, their respective transmitted data signals from the BS (downlink) or their respective received data signals at the BS (uplink) could potentially interfere with one another. The exact degree of this interference depends on the spatial correlation between the channels of the paired UEs, the type of MU-MIMO precoder used at the BS (downlink), and the type of MU-MIMO detector used at the BS (uplink). Therefore, the resulting UE data rates depend on how the frequency resources are shared between the UEs and this in turn influences the respective UE ranks and power allocation. Much of the literature has focused on optimizing these three procedures separately, but this leads to numerous balloon-effects where optimizing one resource reduces the gains achievable by optimizing another. Furthermore, up until the recent advent of 5G NR, MU-MIMO was not commonly used which has also delayed the development of joint solutions for this problem. Practical solutions for the joint allocation of time-frequency resources, power control, and rank selection remain elusive and are much needed for 6G MU-MIMO to succeed.
This paper addresses the problem of joint resource-power allocation and rank selection for a class of linear transceivers. In particular, we consider zero-forcing (ZF) precoders [1] for downlink (DL) transmission and linear minimum mean square error (LMMSE) receivers for uplink (UL) reception. We show that both these problems can be formulated under a common framework. Since the problems involve mixed integer programming for which finding the optimal solution could be computationally intensive, we provide a practically-feasible approach to obtain a reasonably good (even if possibly sub-optimal) solution. To the best of our knowledge, this is the first such work on jointly selecting the three aforementioned features based on the type of precoder or receiver used.
@@
Notation: Boldface upper-case (lower-case) letters denote matrices (vectors). The field of complex numbers and the field of real numbers are respectively denoted by $\mathbb{C}$ and $\mathbb{R}$. For any set ${\cal S}$, $|{\cal S}|$ denotes its cardinality if it is finite and the notation $\mathbf{X}\in{\cal S}^{m\times n}$ denotes that $\mathbf{X}$ is a matrix of size $m\times n$ with each entry taking values from ${\cal S}$. For a matrix $\mathbf{X}$, its transpose and Hermitian transpose are respectively denoted by $\mathbf{X}^{T}$ and $\mathbf{X}^{H}$, Frobenius norm by $\|\mathbf{X}\|$, and its $(i,j)^{th}$ entry by $[\mathbf{X}]_{i,j}$. The block-diagonal matrix with diagonal blocks $\mathbf{D}_{1},\mathbf{D}_{2},\cdots,\mathbf{D}_{n}$ is denoted by $\mathrm{diag}\left[\mathbf{D}_{1},\mathbf{D}_{2},\cdots,\mathbf{D}_{n}\right]$, and the same notation is used for denoting a diagonal matrix with diagonal elements $d_{1},\cdots,d_{n}$. The identity matrix and the null matrix are respectively denoted by $\mathbf{I}$ and $\mathbf{O}$ with their sizes understood from context. Finally, $\mathbb{I}_{\{A\}}$ denotes the indicator function for event $A$; $\mathbb{I}_{\{A\}}=1$ if $A$ occurs, and is 00 otherwise.
@@ II System Model
We consider an OFDM-based multiple-input multiple-output (MIMO) communication system with $n_{F}$ subcarriers. A resource element (RE) corresponds to a subcarrier-OFDM symbol pair in the OFDM grid. A slot consists of $T$ OFDM symbols (usually $T=12$ or $14$) and a PRB consists of 12 subcarriers. Suppose that the BS is equipped with $n_{B}$ antennas and serves a set of $N_{UE}\geq 1$ co-scheduled UEs, with UE $i$ equipped with $n_{U}^{(i)}$ antennas, $i=1,\cdots,N_{UE}$. Each UE $i$ is assigned a rank (also called the number of data layers or spatial streams assigned for that UE) $n_{l}^{(i)}\leq n_{U}^{(i)}$ with $n_{L}\triangleq\sum_{i=1}^{N_{UE}}n_{l}^{(i)}\leq n_{B}$. Let $\mathbf{H}_{i,f,t}\in\mathbb{C}^{n_{B}\times n_{U}^{(i)}}$ denote the channel matrix between UE $i$ and the BS for RE $(f,t)$, where $f$ corresponds to the subcarrier index and $t=1,\cdots,T$, to the OFDM symbol index. While it is assumed that the BS has perfect knowledge of this channel-state information (CSI), in practice, only imperfect estimates are available, either through sounding reference signal (SRS) [8, Sec. 6.4.1.4] transmission on the uplink in a time division duplex (TDD) system or through UE CSI feedback in a frequency division duplex (FDD) system. Moreover, these estimates are only available at a certain granularity (e.g., one channel estimate for every two consecutive PRBs). In this paper, we call this set of contiguous PRBs for which a single channel estimate is available a physical resource block group (PRBG). This can also be a group of PRBs over which the channel is approximately constant. Therefore, resource allocation refers to the assignment of PRBGs to the $N_{UE}$ UEs that are to be served in the slot in context in order to optimize a certain performance metric, with possibly multiple UEs sharing the same PRBG. It is assumed that once a PRBG is assigned to a UE, all the $T$ OFDM symbols are used for communication by the UE in that PRBG. The constellation used for communication between the BS and UE $i$ is a unit energy QAM constellation of size $2^{m_{i}}$ for some $m_{i}=2,4,6,8,\cdots$, and is denoted by ${\cal Q}_{i}$.
@@ II-B Downlink transmission
With ${\cal I}_{f,t}$ and $n_{L,f,t}$ as defined in the previous subsection, let $\mathbf{W}_{i,f,t}\in\mathbb{C}^{n_{B}\times n_{l}^{(i)}}$ denote the part of the precoding matrix associated with UE $i$ if $i\in{\cal I}_{f,t}$. Then, the overall precoding matrix $\mathbf{W}_{f,t}\in\mathbb{C}^{n_{B}\times n_{L,f,t}}$ used by the BS is obtained by stacking the matrices $\{\mathbf{W}_{i,f,t},i\in{\cal I}_{f,t}\}$ column-wise, with each column of $\mathbf{W}_{f,t}$ having unit norm. Further, let $a_{k,i,j,f,t}\triangleq|\left[\mathbf{W}_{i,f,t}\right]_{k,j}|^{2}$, $k=1\cdots,n_{B},j=1,\cdots,n_{l}^{(i)}$. The BS transmits $\mathbf{x}_{i,f,t}\triangleq\left[x_{i,1,f,t},\cdots,x_{i,n_{l}^{(i)},f,t}\right]^{T}\in{\cal Q}_{i}^{n_{l}^{(i)}\times 1}$ on RE $(f,t)$ to UE $i$ if $i\in{\cal I}_{f,t}$, and uses a transmit power $p_{i,j,f,t}>0$ for the $j^{th}$ layer of UE $i$ ($p_{i,j,f,t}=0$ if $i\notin{\cal I}_{f,t}$). Let $\mathbf{D}_{i,f,t}\triangleq\mathrm{diag}\left[\sqrt{p_{i,1,f,t}},\cdots,\sqrt{p_{i,n_{l}^{(i)},f,t}}\right]$. With a maximum BS transmit power of $P_{B,max}$ and a per-antenna power constraint (PAPC) of $P_{ant}=P_{B,max}/n_{B}$, it follows that $\forall t=1,\cdots,T$,
$\sum_{k=1}^{n_{B}}\sum_{i\in{\cal I}_{f,t}}\sum_{j=1}^{n_{l}^{(i)}}\sum_{f=1}^{n_{F}}a_{k,i,j,f,t}p_{i,j,f,t}$$\leq$ $P_{B,max},$ (6)
$\sum_{i\in{\cal I}_{f,t}}\sum_{j=1}^{n_{l}^{(i)}}\sum_{f=1}^{n_{F}}a_{k,i,j,f,t}p_{i,j,f,t}$$\leq$ $P_{ant},$ (7)
$\forall k=1,\cdots,n_{B}$ in (7). In this paper, we consider the class of precoders called ZF precoders [1] that satisfy, for any $j,i\in{\cal I}_{f,t}$, $\mathbf{H}_{j,f,t}^{T}\mathbf{W}_{i,f,t}=\mathbf{O}$ when $j\neq i$, and $\mathbf{H}_{i,f,t}^{T}\mathbf{W}_{i,f,t}=\mathbf{U}_{i,f,t}\mathbf{\Lambda}_{i,f,t}$ for some matrix $\mathbf{U}_{i,f,t}\in\mathbb{C}^{n_{U}^{(i)}\times n_{l}^{(i)}}$ satisfying $\mathbf{U}_{i,f,t}^{H}\mathbf{U}_{i,f,t}=\mathbf{I}$ and $\mathbf{\Lambda}_{i,f,t}=\mathrm{diag}\left[\sqrt{\lambda_{i,1,f,t}},\cdots,\sqrt{\lambda_{i,n_{l}^{(i)},f,t}}\right]$ with positive real-valued diagonal elements. With this, the signal model for the received signal at UE $i$ on RE $(f,t)$ is
$\mathbf{y}_{i,f,t}=\mathbf{H}_{i,f,t}^{T}\mathbf{W}_{i,f,t}\mathbf{D}_{i,f,t}\mathbf{x}_{i,f,t}+\mathbf{n}_{i,f,t}$ (8)
@@ II-C Expected UE Rates
Modern communication systems use bit-interleaved coded modulation (BICM) [10] where, due to the usage of scrambling and interleaving, the transmitted bits in an RE are approximately independent and identically distributed. Let ${\cal G}_{i}\triangleq\{(f,t)|i\in{\cal I}_{f,t}\}$ denote the indices of the RE allocated for communication between the BS and UE $i$. In practical communication systems, there is finite number of modulation and coding schemes (MCSs). For example, 5G NR Table 2 supports up to $27$ MCS levels [8, Table 5.1.3.1-2], with the lowest level using $4$-QAM and a rate-$\frac{120}{1024}$ low-density parity-check (LDPC) code with spectral efficiency (SE) given by $\left(120/1024\right)\mathop{\mathrm{log}_{2}}\left(4\right)=0.234$, and the highest level using $256$-QAM and a rate-$\frac{948}{1024}$ LDPC code with SE $\left(948/1024\right)\mathop{\mathrm{log}_{2}}\left(256\right)=7.40$. We denote the minimum MCS SE by $r_{min}$ and the maximum SE possible by $r_{max}=\mathop{\mathrm{log}_{2}}\left(|Q_{max}|\right)$, respectively, where $Q_{max}$ is the highest-order constellation used in the system. The achievable rate $R_{UE,i}$ (in bits per slot) for UE $i$ is upper-bounded as
@@
where $\rho_{i,j,f,t,i}$ is the post-equalization SINR for the $j^{th}$ received symbol of UE $i$ in the RE indexed by $(f,t)$, and $I_{Q}(x)$ refers to the mutual-information (MI) between the input and the output in an AWGN channel for constellation $Q$ and an SNR of $x$. Since there is a minimum MCS SE $r_{min}$, reliable communication is not guaranteed if the expression in (11) is lower than $\left|{\cal G}_{i}\right|n_{l}^{(i)}r_{min}$. Therefore, the achievable rate for UE $i$ is
$R_{UE,i}=x\mathbb{I}_{\left\{x>n_{l}^{(i)}|{\cal G}_{i}|r_{min}\right\}}$ (12)
where $x$ where is given by the R.H.S. of (11).